# Optical Properties, Quasinormal Modes and Greybody factors of deformed AdS-Schwarzschild black holes

### v1

@@ VI Greybody Factors
$\begin{array}[]{ccll}\Psi&=&e^{-i\omega r_{*}}+Re^{i\omega r_{*}},&r_{*}\rightarrow+\infty,\\ \Psi&=&Te^{-i\omega r_{*}},&r_{*}\rightarrow-\infty,\\ \end{array}$ (27)
@@
We have shown the variations of greybody factors with respect to the model parameters in Fig.s 13, 14 and 15.
@@
The two panels of Fig. 14 focus on the effect of varying $\alpha$, with $l=2$ and $\beta=0.2$ held constant. As $\alpha$ increases, the onset of significant transmission shifts to higher frequencies, indicating that the deformation of the black hole spacetime increases the potential barrier. This behaviour highlights the strong influence of the deformation parameter on the greybody factors, suggesting that modifications to black hole geometry can have a noticeable effect on the transmission characteristics of perturbations.
@@ VII Emission Rate
As shown in Fig. 16, an increase in the parameter $\alpha$, causes the maximum value of the emission rate to decrease and occur at a lower frequency. Also, an increase in the parameter $\beta$, leads to a decrease in the emission rate maximum value and causes the maximum occurring at a higher frequency value.
VIII Concluding Remarks
@@
In this work, we explored several properties of the deformed black hole spacetime such as the temperature, photon and shadow radii, QNMs, greybody factors and emission rate, mainly focusing on the effects of the deformation parameter $\alpha$ and control parameter $\beta$ on both scalar and electromagnetic perturbations. The investigation of shadow radius as a function of parameters $\alpha,\,\beta$ and $\Lambda$, leads to an upper and lower limit of parameters in which the shadow radius is allowed as a result of the observed data of Sgr A. Our analysis of QNMs revealed that an increase in $\alpha$ significantly enhances the oscillation frequency and damping rate of ringdown GWs while increasing $\beta$ leads to a non-linear behaviour in both frequency and damping rates. Specifically, the damping rate initially increases with $\beta$ before reversing its trend beyond $\beta=0.65$. For electromagnetic perturbations, we observed a similar pattern, though with slightly lower oscillation frequencies and damping rates compared to scalar perturbations. These variations underscore the sensitivity of black hole dynamics to these parameters, which are crucial for understanding astrophysical processes like black hole mergers and gravitational wave emissions.
@@
Additionally, we examined the greybody factors for scalar and electromagnetic perturbations, with a focus on the role of angular momentum $l$, $\alpha$, and $\beta$. Our results indicate that both $l$ and $\alpha$ play a dominant role in shaping the transmission of perturbations through the black hole’s potential barrier. Higher angular momentum leads to stronger scattering while increasing $\alpha$ shifts the onset of significant transmission to higher frequencies. In contrast, $\beta$ has a more subtle impact, fine-tuning the transmission profiles without causing major shifts in the greybody factors.
@@ Acknowledgments
DJG acknowledges the contribution of the COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”. Also, the research was partially supported by the Long-Term Conceptual Development of a University of Hradec Králové for 2023, issued by the Ministry of Education, Youth, and Sports of the Czech Republic.

### v2

@@ VI Greybody Factors
[…]
@@
We have shown the variations of greybody factors with respect to the model parameters in Fig.s 20, 21 and 22.
@@
The two panels of Fig. 21 focus on the effect of varying $\alpha$, with $l=2$ and $\beta=0.2$ held constant. As $\alpha$ increases, the onset of significant transmission shifts to higher frequencies, indicating that the deformation of the black hole spacetime increases the potential barrier. This behaviour highlights the strong influence of the deformation parameter on the greybody factors, suggesting that modifications to black hole geometry can have a noticeable effect on the transmission characteristics of perturbations. To understand the variation of the greybody factors from a standard black hole in the absence of deformation, we have shown a curve with $\alpha=\beta=0.$ It is clearly evident that the presence of deformation increases the potential barrier.
@@ VII Emission Rate
As shown in Fig. 23, an increase in the parameter $\alpha$, causes a decrease in the emission rate of the black hole. As a result, the lifetime of the black hole increases in comparison to that of a black hole without deformation. On the other hand, an increase in the parameter $\beta$, leads to an increase in the emission rate and as a result, the life span of the black hole decreases.
VIII Topological Characteristics
In this section, the stability of the photon sphere and the classification of black holes based on thermodynamic potentials will be investigated using a topological approach.
VIII.1 Topological Charge of Photon Sphere
Photon sphere of the black hole of Eq. (4) is explored in Sec. IV. In order to study the stability or instability of the photon sphere, a topological approach is employed. To this end, a potential of form wei2020topological ; sadeghi2024role
$H(r,\,\theta)=\frac{1}{\sin\theta}\frac{\mathcal{F}(r)}{r^{2}}$ (37)
is defined, which can be represented in a vector space using wei2020topological
$\phi^{H}=\left(\frac{\partial}{\partial r}H(r,\,\theta),\,\frac{\partial}{\partial\theta}H(r,\,\theta)\right)$ (38)
which can be normalized as
[…]
Each photon sphere in vector space is represented as a zero point, to which a topological charge can be assigned. To study the stability of the zero point, the sign of the topological charge arising from the rotation of vectors around this point is examined as zhu2024topological
$w_{i}=\frac{1}{2\pi}\oint_{c}d(\arctan\frac{n_{\theta}}{n_{r}}).$ (40)
For this purpose, a closed contour is taken around the zero point $r_{i}$, and a change of variables in the form of
$r=a\cos\vartheta+r_{i},\;\;\;\;\theta=b\sin\vartheta+\frac{\pi}{2},$ (41)
is utilized to calculate the topological charge of the zero point. Fig. 24 illustrates the vector space of potential $H(r,\,\theta)$ for the case $M=1$, $\alpha=0.20$, $\beta=0.20$, and $\Lambda=-0.002$, and as shown, a zero point is located at $r_{c}=2.88232$.
Refer to caption    Refer to caption
Figure 24: Assuming $M=1$, $\alpha=0.20$, $\beta=0.20$, and $\Lambda=-0.002$. Left panel: Vector space of potential $H(r,\,\theta)$, zero point is located at $r_{c}=2.88232$. Right panel: Variation of $\Delta\Omega$ versus $\vartheta$, considering $a=b=0.3$.
Considering $a=b=0.3$, the topological charge of the photon sphere located at $r_{c}$ equals -1, indicating an unstable photon sphere. This instability is consistent with our studies in Sec. IV, as this unstable photon sphere leads to the formation of the shadow.
VIII.2 Topological Charge of Thermodynamic Potentials
Similar to the previous section, the thermodynamic potentials of the black hole can be studied using a topological approach. Hawking temperature of the black hole was calculated in Eq. (8). This temperature can be expressed in isobaric form as
$T_{i}=\frac{1}{2\sqrt{\pi S}}-\frac{\sqrt{\pi}\alpha S}{\left(\sqrt{\pi}\beta+\sqrt{S}\right)^{5}}$ (42)
The curve of isobaric temperature variations as a function of $\sqrt{S}$ is presented in Fig. 25 for the case $\alpha=0.40$ and $\beta=0.18$.
Refer to captionRefer to captionRefer to caption
Figure 25: Assuming $\alpha=0.40$ and $\beta=0.18$. Left panel: Isobaric curve of Hawking temperature. Middle Panel: Vector space of potential $\Phi$. Right panel: Variation of $\Delta\Omega$ versus $\vartheta$, considering $a=b=0.1$.
The Hawking temperature–dependent potential is defined as wei2022topology ; mehmood2023thermodynamic
$\Phi=\frac{1}{\sin\theta}T_{i}.$ (43)
The vectors of the above potential can be represented using wei2022topology
$\phi^{\Phi}=(\frac{\partial}{\partial r_{h}}\Phi,\,\frac{\partial}{\partial\theta}\Phi),$ (44)
This vector space for cases $\alpha=0.40$ and $\beta=0.18$ is depicted in Fig. 25. This vector space has two zero points enclosed by contours $c_{1}$ and $c_{2}$. Considering $a=b=0.1$, their topological charges are +1 and 1-1, respectively, indicating a novel phase transition and a conventional phase transition wei2022topology ; yerra2022topology .
Generalized free energy of a black hole is another quantity which can be analyzed thermodynamically and is defined as wei2022black ; wu2025novel
$F=M(r_{h})-\frac{S}{\tau}$ (45)
that $M(r_{h})$ is black hole’s mass as a function of horizon radius, $\tau$ is the inverse of black hole temperature, and on the horizon becomes $\tau\big{|}_{r=r_{h}}=T_{H}$, and $S$ indicates entropy of the black hole and is calculated as
$S=\int\frac{dM(r_{h})}{T}dr_{h}=\pi r_{h}^{2}$ (46)
Thus, off–shell generalized free energy of the black hole of form Eq. (4) reads
$F=\frac{r_{h}}{2}+\frac{\alpha\left(\beta^{2}+3r_{h}^{2}+3\beta r_{h}\right)}{(\beta+r_{h})^{3}}-\frac{\Lambda r_{h}^{3}}{6}-\frac{\pi r_{h}^{2}}{\tau}$ (47)
and vectors of the above potential are defined as wei2022black ; fang2023revisiting
$\phi^{F}=(\frac{\partial}{\partial r_{h}}F,\,-\cot\theta\csc\theta).$ (48)
Zero point of potential field $F$ is located at $\dfrac{\partial}{\partial r_{h}}F=0$. Thus parameter $\tau$ at the zero point can be found as function of $r_{h}$ as
$\tau=\frac{4\pi r_{h}(\beta+r_{h})^{4}}{\beta^{4}+r_{h}^{4}+4\beta r_{h}^{3}-\alpha r_{h}^{2}+6\beta^{2}r_{h}^{2}-\Lambda r_{h}^{2}(\beta+r_{h})^{4}+4\beta^{3}r_{h}}$ (49)
Fig. 26 illustrates $r_{h}-\tau$ curve for the given set of initial values $\alpha=0.20$, $\beta=0.20$, and $P=0.0022$. As shown, As can be seen, this curve changes direction at (0.45782, 9.94262)(0.45782,\,9.94262) and (4.18812, 26.8591)(4.18812,\,26.8591).
Refer to captionRefer to captionRefer to caption
Figure 26: Assuming $\alpha=0.20$, $\beta=0.20$, and $P=0.0022$. Middle panel: Vector space of potential $\mathcal{F}$, when $\tau=26$. Right panel: Variation of $\Delta\Omega$ versus $\vartheta$, considering $a=b=0.3$.
By choosing $\tau=26$, the vector space of the potential $F$ is illustrated in Fig. 26. This vector space includes three zero points at $r_{h_{1}}=0.23014$, $r_{h_{2}}=3.21936$, and $r_{h_{3}}=5.43388$, with their topological charges equal to +1+1, 1-1, and +1+1, respectively, where $a=b=0.3$. For the given set of initial values, the whole topological charge is +1+1. Therefore, this black hole is classified within a class analogous to RN–AdS black hole wei2022black .
IX Concluding Remarks
@@
In this work, we explored several properties of the deformed black hole spacetime, such as the temperature, photon and shadow radii, QNMs, greybody factors and emission rate, mainly focusing on the effects of the deformation parameter $\alpha$ and control parameter $\beta$ on both scalar and electromagnetic perturbations. The investigation of shadow radius as a function of parameters $\alpha,\,\beta$ and $\Lambda$, leads to an upper and lower limit of parameters in which the shadow radius is allowed as a result of the observed data of Sgr A. Our analysis of QNMs revealed that an increase in $\alpha$ significantly enhances the oscillation frequency and damping rate of ringdown GWs while increasing $\beta$ leads to a non-linear behaviour in both frequency and damping rates. Specifically, the damping rate initially increases with $\beta$ before reversing its trend beyond $\beta=0.65$. For electromagnetic perturbations, we observed a similar pattern, though with slightly lower oscillation frequencies and damping rates compared to scalar perturbations. These variations underscore the sensitivity of black hole dynamics to these parameters, which are crucial for understanding astrophysical processes like black hole mergers and gravitational wave emissions. This investigation ascertains that deformation parameters $\alpha$ and $\beta$ play distinct roles in shaping the black hole geometry and, consequently, influence the behaviour of QNM in different ways. The parameter $\alpha$ governs the overall strength of the deformation introduced through the energy density profile, effectively modifying the depth and curvature of the effective potential experienced by perturbations. As $\alpha$ increases, the potential barrier becomes steeper and more localized, leading to higher oscillation frequencies ($\omega_{R}$) and faster damping rates ($|\omega_{I}|$), as is typical in systems with sharper confinement. This is consistent with the intuitive expectation that stronger spacetime curvature near the peak of the potential enhances the restoring force on perturbations, thus increasing the QNM frequency.
On the other hand, $\beta$ controls the regularization scale near the black hole core and has a more subtle influence. Its primary effect is to alter the inner structure of the geometry without significantly modifying the asymptotic form. As shown in the results, the behaviour of $\omega_{I}$ with $\beta$ is non-monotonic, displaying a reversal near a critical value (e.g., $\beta\approx 0.65$), which can be interpreted as a transition in the internal structure of the effective potential. For small $\beta$, the regularizing effect is weak, and the potential resembles that of a singular black hole; increasing $\beta$ initially strengthens the damping, but beyond a threshold, the regularization dominates and leads to a broadening of the potential barrier, reducing the decay rate. This non-linear behaviour reflects the competition between enhanced regularity and the geometric sharpness of the potential barrier, and it may have interesting implications for the stability and ringdown profiles of regular or exotic compact objects.
While our study reveals that the deformation parameters $\alpha$ and $\beta$ introduce significant modifications to the standard QNM spectrum -particularly enhancing the oscillation frequency and altering the damping behavior—translating these theoretical results into observational constraints presents a challenge with current detection capabilities. Ground-based gravitational wave detectors such as LIGO and Virgo are primarily sensitive to QNM frequencies within the range $f\in[12\,\text{Hz},1.2\,\text{kHz}]$. By converting the dimensionless QNM frequencies into physical units using the relations derived in Ref. Ferrari2008 , we estimate the mass ranges of black holes whose ringdown signals could be detected within this sensitivity window. Our analysis shows that, for both types of perturbations (scalar, and electromagnetic), stellar-mass black holes with masses between approximately $8M_{\odot}$ and $1100M_{\odot}$ are within the detectable range for LIGO-Virgo. However, for supermassive black holes such as Sagittarius A ($M\sim 3.7\times 10^{6}M_{\odot}$), the corresponding QNM frequencies fall well below the LIGO-Virgo band but lie comfortably within the future LISA frequency range ($0.1\,\text{mHz}$ to $1\,\text{Hz}$). Therefore, although current ground-based detectors may not provide stringent constraints on the deformation parameters, space-based observatories like LISA offer promising avenues for probing such modifications through the precise measurement of QNMs from massive black holes. In this context, our results provide a theoretical basis for identifying potential signatures of deviations from general relativity, such as those induced by deformation parameters, in upcoming gravitational wave observations.
@@
Additionally, we examined the greybody factors for scalar and electromagnetic perturbations, with a focus on the role of angular momentum $l$, $\alpha$, and $\beta$. Our results indicate that both $l$ and $\alpha$ play a dominant role in shaping the transmission of perturbations through the black hole’s potential barrier. Higher angular momentum leads to stronger scattering, while increasing $\alpha$ shifts the onset of significant transmission to higher frequencies. In contrast, $\beta$ has a more subtle impact, fine-tuning the transmission profiles without causing major shifts in the greybody factors.
We have also investigated the topological characteristics of the black hole. We found that this black hole can be classified within a class analogous to RN–AdS black hole from this perspective.
@@ Acknowledgments
The authors would like to thank the referees for their constructive comments. DJG acknowledges the contribution of the COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”. F. S. and H. H. are grateful to the Excellence project FoS UHK 2203/2025-2026 for the financial support. Also, the research of H.H. was supported by the Q-CAYLE project, funded by the European Union-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon (PRTRC17.11), and also by project PID2023-148409NB-I00, funded by MICIU/AEI/10.13039/501100011033. Financial support of the Department of Education of the Junta de Castilla y Leon and FEDER Funds is also gratefully acknowledged (Reference: CLU-2023-1-05). The authors would like to thank Prof. R. A. Konoplya for his useful discussion.